# Improved Bug Localization with AI Agents Leveraging Hypothesis and Dynamic Cognition

### v1

@@ 5. Related Work
5.1. IR and Bug Localization
Traditional bug localization methods fall into two categories: spectra-based and Information Retrieval (IR)-based (Wang et al., 2015a). Spectra-based approaches rely on execution traces and test cases, making them resource-intensive and often impractical (Moreno et al., 2014; Wang et al., 2015b). On the other hand, IR-based methods treat bug localization as a retrieval task and rely on the textual similarity between bug reports and source code to identify buggy documents.
@@ 5.1. IR and Bug Localization
Since IR performance largely depends on query quality, some researchers focused on query reformulation—refining bug report queries by adding or removing terms (Mills et al., 2020; Rahman et al., 2021). Refoqus (Haiduc et al., 2013) employed machine learning to suggest expansion or reduction strategies, while graph-based and genetic methods (Rahman and Roy, 2018; Rahman et al., 2021) optimized queries by analyzing term relationships. Relevance feedback techniques, such as Rocchio’s algorithm (Haiduc et al., 2013) and Spatial Code Proximity (SCP) (Sisman and Kak, 2013), further reformulated queries based on statistical and co-occurrence patterns from initial search results.
@@
5.2. Deep Learning and Bug Localization
Advances in deep learning and Large Language Models (LLMs) have improved the understanding of source code and natural language, motivating their application to bug localization. Existing approaches can be broadly categorized into training-based models and prompt-based instruction-following methods.
Among training-based techniques, DNNLOC (Lam et al., 2017a), a seminal work, trains on positive and negative (bug report, source file) pairs and incorporates multiple signals (e.g., rVSM similarity (Kim et al., 2013), class-name similarity, bug-fix recency), with a second network combining them for ranking. However, its dependence on bug-fixing recency can limit practical use. FBL-BERT (Ciborowska and Damevski, 2022) trains on bug-report and changeset pairs and follows ColBERT (Khattab and Zaharia, 2020)-style late interaction for ranking, matching bug-report tokens to their most similar tokens in each changeset to identify buggy code. However, its reliance on changesets limits applicability in rapidly evolving projects. BLAZE (Chakraborty et al., 2025) trains a GPT-based model using hard-example contrastive learning and dynamic chunking to improve alignment between bug reports and source code, enabling effective retrieval of buggy documents. Other approaches have employed convolutional neural networks to learn representations for bug localization (Huo and Li, 2017; Ali et al., 2023). However, these techniques can lack generalizability and require retraining, limiting scalability in large, evolving codebases.
@@ 6. Threats to Validity
Threats to internal validity, which involve potential experimental errors or biases, primarily arise from replicating existing baselines. To address this, we relied on replication packages released by the original authors (e.g., BLIZZARD, BRaIn, Agentless, LocAgent (Rahman, ; Xia et al., 2024; Samir and Rahman, 2025a; Chen et al., 2025)) and from Bench4BL (Lee et al., 2018) (BLUiR (Saha et al., 2013)). Since Indri is now deprecated, we replaced it with Lucene in our BLUiR replication. To reduce bias, we tested on two datasets and found minimal differences from the baselines.
Threats to external validity pertain to the generalizability of our findings. Although CogniGent was evaluated solely on Java projects, the underlying language models (e.g., LLaMA (AI, )) are inherently adaptable to multiple programming languages, which helps reduce this limitation.
@@ 7. Conclusion
Software bugs undermine system reliability and security, leading to significant financial losses (Zahn et al., 2024; Wee, 2024). To address these long-standing challenges, we introduce CogniGent, a technique designed to overcome the limitations of IR and AI-based bug localization by enabling causal reasoning, dependency-aware navigation, and context-managed code exploration. CogniGent applies human-inspired Dynamic Cognitive Debugging within an LLM-driven agentic workflow to formulate hypotheses, traverse interconnected components, and trace fault propagation to identify underlying causes. To evaluate CogniGent’s performance, we use three widely adopted metrics in bug localization—Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K—and observed substantial improvements: 23.33% and 38.57% gains in MAP at the document and method levels, respectively, and 25.14% and 53.74% gains in MRR across the same levels of granularity.
Building on this, we plan to capture finer-grained structural relationships and extend our agentic workflow to better assist developers in locating and resolving bugs.

### v2

@@ 5. Related Work
IR and Bug Localization. Traditional bug localization methods fall into two categories: spectra-based and Information Retrieval (IR)-based (Wang et al., 2015a). Spectra-based approaches rely on execution traces and test cases, making them resource-intensive and often impractical (Wang et al., 2015b). On the other hand, IR-based methods treat bug localization as a retrieval task and rely on the textual similarity between bug reports and source code to identify buggy documents.
@@
Since IR performance largely depends on query quality, some researchers focused on query reformulation—refining bug report queries by adding or removing terms (Mills et al., 2020; Rahman et al., 2021). Refoqus (Haiduc et al., 2013) employed machine learning to suggest expansion or reduction strategies, while graph-based and genetic methods (Rahman and Roy, 2018; Rahman et al., 2021) optimized queries by analyzing term relationships. Relevance feedback techniques, such as Rocchio’s algorithm (Haiduc et al., 2013) and Spatial Code Proximity (SCP) (Sisman and Kak, 2013), further reformulate queries based on statistical and co-occurrence patterns from initial search results.
@@
Deep Learning and Bug Localization. Advances in deep learning and Large Language Models (LLMs) have improved the understanding of source code and natural language, motivating their application to bug localization. Existing approaches can be broadly categorized into training-based models and prompt-based instruction-following methods.
Among training-based techniques, DNNLOC (Lam et al., 2017a), a seminal work, trains on positive and negative (bug report, source file) pairs and incorporates multiple signals (e.g., rVSM similarity (Kim et al., 2013), class-name similarity, bug-fix recency), with a second network combining them for ranking. However, its dependence on bug-fixing recency can limit practical use. FBL-BERT (Ciborowska and Damevski, 2022) trains on bug-report and changeset pairs and follows ColBERT (Khattab and Zaharia, 2020)-style late interaction for ranking, matching bug-report tokens to their most similar tokens in each changeset to identify buggy code. However, its reliance on changesets limits its applicability in rapidly evolving projects. BLAZE (Chakraborty et al., 2025) trains a GPT-based model using hard-example contrastive learning and dynamic chunking to improve alignment between bug reports and source code, enabling effective retrieval of buggy documents. Other approaches have employed convolutional neural networks to learn representations for bug localization (Huo and Li, 2017; Ali et al., 2023). However, these techniques can lack generalizability and require retraining, limiting scalability in large, evolving codebases.
@@ 6. Threats to Validity
Threats to internal validity, which involve potential experimental errors or biases, primarily arise from replicating existing baselines. To address this, we relied on replication packages released by the original authors (e.g., BLIZZARD, BRaIn, Agentless, LocAgent (Rahman, ; Xia et al., 2024; Samir and Rahman, 2025a; Chen et al., 2025)) and from Bench4BL (Lee et al., 2018) (BLUiR (Saha et al., 2013)). Since Indri (Strohman et al., 2005) is now deprecated, we replaced it with Lucene in our BLUiR replication. To reduce bias, we tested on two datasets and found minimal differences from the baselines.
Threats to external validity pertain to the generalizability of our findings. Although CogniGent was evaluated solely on Java projects, the underlying language models (e.g., LLaMA (AI, )) are inherently adaptable to multiple programming languages, which helps mitigate this limitation.
@@
Finally, we opted for locally deployable, open-weight models (e.g., LLaMA (AI, )) for our experiments, as reproducibility with proprietary models (e.g., GPT-4 (OpenAI, 2023)) might be affected by cost barriers. Additionally, we release the replication package (15), which mitigates the risks associated with experimental configurations (e.g., temperature = 0.5), but results may vary due to LLM non-determinism (Song et al., 2025).
@@ 7. Conclusion
Software bugs undermine system reliability and security, leading to significant financial losses (Zahn et al., 2024; Wee, 2024). To address these long-standing challenges, we introduce CogniGent, a technique designed to overcome the limitations of IR and AI-based bug localization by enabling causal reasoning, dependency-aware navigation, and context-managed code exploration. CogniGent applies human-inspired Dynamic Cognitive Debugging within an LLM-driven agentic workflow to formulate hypotheses, traverse interconnected components, and trace fault propagation to identify underlying causes. To evaluate CogniGent’s performance, we use three widely adopted metrics in bug localization–Mean Average Precision (MAP), Mean Reciprocal Rank (MRR), and HIT@K—and observed substantial improvements: 23.33% and 38.57% gains in MAP at the document and method levels, respectively, and 25.14% and 53.74% gains in MRR across the same levels of granularity. Building on these results, we plan to capture finer-grained structural relationships and extend our agentic workflow to further support bug resolution.
Acknowledgements.
This work was supported by the Natural Sciences and Engineering Research Council of Canada (Discovery Grant RGPIN‑03236)